OpenStudy (i<3mybaby):

The ratio of the measures of two supplementary angles is 7:3. Find the measures of the angles. A. 126° and 54° B. 72° and 18° C. 159° and 21° D. 69° and 21°

sammixboo (sammixboo):

I believe it is A... If you look at this problem it explains it more... http://openstudy.com/study#/updates/504ecf8ee4b05be998159054

OpenStudy (anonymous):

supplementary angles add up to 180 so it can't be the second or the last ones and 126 + 54 = 180 and 159 + 21 = 180 so what you want to do is find 126/ 7 and 54/7 (do the same thing for 159 and 21) and if they end up being the same, then you have your answer

OpenStudy (i<3mybaby):

@jadeybells... wait isnt it 126/7 and 54/3?

OpenStudy (anonymous):

that's what I meant /).< sorry

OpenStudy (i<3mybaby):

it's ok so that one is 18 and 18 and 159/7 =22.7. but 21/3=7 so it has to be a

OpenStudy (anonymous):

Yep :) Like Sammi said, as well

sammixboo (sammixboo):

*dances*

OpenStudy (i<3mybaby):

thanks